OpenStudy (anonymous):

What is the slope of the line if its : (2,6) and (1,3) A. 2/1 B. 3/2 C. 3/1 D. 1/3

TheSmartOne (thesmartone):

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

TheSmartOne (thesmartone):

Also, \(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @XoXoTheWeeknd @~ :D

OpenStudy (anonymous):

thanks for the welcoming :) but i dont understand slopes at all...im kind of new to this

TheSmartOne (thesmartone):

What are the 2 y-values from the two points? :)

OpenStudy (anonymous):

6 and 3? @TheSmartOne

OpenStudy (braydonlevi99):

Yes you're right Xoxo

OpenStudy (anonymous):

thnks :) what would be the next step? im taking notes for review

OpenStudy (anonymous):

Use http://www.geogebra.org/ and make a line, find rise over run, and there is your slope! This is a very helpful website! let me know if you need anything else.

TheSmartOne (thesmartone):

So 6 - 3 = 3 That's the numerator What are the 2 x-values?

OpenStudy (anonymous):

2 and 1 @TheSmartOne

TheSmartOne (thesmartone):

and 2 - 1 = 1 What is 3/1 = ?

TheSmartOne (thesmartone):

since 3 is our numerator, and 1 is our denominator we get the fraction as 3/1 which is the slope :)

OpenStudy (anonymous):

thanks that was very easier then i thought :)

TheSmartOne (thesmartone):

good to hear :)

TheSmartOne (thesmartone):

If your question has been answered, you can press the "Close button"

TheSmartOne (thesmartone):

http://prntscr.com/9dvzyb